# MARS-M: When Variance Reduction Meets Matrices

### v2

@@ Appendix A Hyper-parameter Settings
For training parameters, for experiments with either OpenWebText or FineWeb-Edu 100B datasets, we do a grid search over learning rates between $\{5e-4,1e-3,3e-3,5e-3,6e-3,1e-2\}$ for Moonlight optimizer. We just apply the same learning rate for MARS-M optimizer. Since Moonlight and MARS-M optimizers are only designed for matrix parameters, we optimize the vector-like parameters and embeddings with AdamW with the same learning rate. We list the architectural hyperparameters for GPT-2 models with 125M (small), 355M (medium), 770M (large) and 1.5B (XL) parameters in Table 2. Moreover, the general hyperparameters used across all experiments are summarized in Table 3, while Table 4 display the training hyperparameters for models with different sizes.
Table 2: Architecture hyperparameters for GPT-2 series models (Radford et al., 2019).
@@
Table 3: General hyper-parameters for the experiments.
Hyper-parameter Value
Steps 100,000
Batch size in total 480
Context length 1024
Gradient clipping threshold 1.0
Dropout 0.0
Learning rate schedule Cosine
Warm-up steps 2000
Base seed 5000
Table 4: Hyper-parameters for GPT-2 experiments for different datasets. We use $\beta=0.95$ for the optimizers.
@@ Appendix B Additional Experiments
We just display the curves for training and validation losses for small, medium and large models on the FineWeb-Edu 100B and OpenWebText datasets in Figures 6-8 and Figures 9-11, respectively. And the 0-shot and 2-shot evaluation results for these models in Tables 5-11 and Tables 12-17, respectively. The results of AdamW are taken from Yuan et al. (2025). According to these figures and tables, it can be observed that MARS-M can achieve better performances than AdamW and Moonlight optimizers in most of these cases.
@@
Figure 7: The training and validation loss of medium-size models (355M) trained with different optimizers on the FineWeb-Edu 100B dataset.
Figure 8: The training and validation loss of large-size models (770M) trained with different optimizers on the FineWeb-Edu 100B dataset.
Figure 9: The training and validation loss of small-size models (125M) trained with different optimizers on the OpenWebText dataset.
Figure 10: The training and validation loss of medium-size models (355M) trained with different optimizers on the OpenWebText dataset.
Figure 11: The training and validation loss of large-size models (770M) trained with different optimizers on the OpenWebText dataset.
@@
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
AdamW 33.53 66.46 45.02 25.32 35.40 69.10 86.70 55.80 52.17
@@
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
AdamW 38.05 70.29 50.30 26.87 38.20 70.46 92.10 55.80 55.26
@@
We have the following lemma from Yuan et al. (2025):
Lemma C.1 (Yuan et al. 2025).
@@
and
$G_{t+1}:=(1-\beta_{t+1})\mathbb{E}\bigg\langle\bm{\Delta}_{t},\nabla f(\mathbf{X}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{X}_{t+1})\bigg\rangle+\beta_{t+1}\mathbb{E}\bigg\langle\nabla F(\mathbf{X}_{t+})-\nabla F(\mathbf{X}_{t}),F(\mathbf{X}_{t})-\mathbf{M}_{t}\bigg\rangle.$
We also need the following lemmas, which can be seen as an extension of Lemma C.3 in Yuan et al. (2025).
@@
Here $\mathbf{O}_{t}=\mathbf{U}_{t}\mathbf{V}_{t}^{\top}$ is defined in (4.1). According to the definition of $\mathbf{O}_{t}$, we have:
@@
$\leq\frac{\rho_{t}}{16L^{2}}\cdot\Big(\frac{\beta_{t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big)\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\rho_{t}}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\frac{\rho_{t}c^{2}\eta_{t}\sigma^{2}}{8L^{2}}-\frac{\rho_{t}}{16L^{2}\eta_{t}}\widetilde{M}_{t+1},$ (C.4)
@@
From Lemma C.3, we know that $\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\leq\eta_{t}^{1/2}$. Since $\rho_{t}^{2}=32L^{2}\eta_{t}$, we obtain [^4] As long as $\eta_{t}\leq 1/2$, $1\geq\beta_{t+1}=1-2\eta_{t}>=0$ is satisfied.
$\frac{\rho_{t}}{16L^{2}}\cdot\Big(\frac{\beta_{t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big)\leq\frac{\rho_{t}}{16L^{2}}\cdot\Big(\eta_{t}^{1/2}-c\Big)\leq\frac{\rho_{t}}{16L^{2}}\cdot\Big(1-c\Big)=-2\eta_{t}\rho_{t}^{-1}.$ (C.5)
@@
$I_{2}\leq-\frac{2\eta_{t}}{\rho_{t}}\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\rho_{t}}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\frac{\rho_{t}c^{2}\eta_{t}\sigma^{2}}{8L^{2}}-\frac{\rho_{t}}{16L^{2}\eta_{t}}\widetilde{M}_{t+1}.$ (C.6)
@@
$\Phi_{t+1}-\Phi_{t}$ $\leq-\frac{\eta_{t}}{\rho_{t}}\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}-\eta_{t}\mathbb{E}\|\mathbf{M}_{t}\|_{F}+\frac{\rho_{t}}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\frac{\rho_{t}c^{2}\eta_{t}\sigma^{2}}{8L^{2}}$
@@
$\leq\Phi_{1}-\Phi_{T+1}+\frac{\sigma^{2}c^{2}}{8L^{2}}\sum_{t=1}^{T}\rho_{t}\eta_{t}+\sum_{t=1}^{T}\big(\frac{3\rho_{t}\eta_{t}n}{8}+\frac{L\eta_{t}^{2}n}{2}\big)-\sum_{t=1}^{T}\frac{\sqrt{2}}{4L\sqrt{\eta_{t}}}\widetilde{M}_{t+1}$

### v3

@@ Appendix A Hyper-parameter Settings
We list the architectural hyperparameters for GPT-2 models with 125M (small), 355M (medium), 770M (large) and 1.5B (XL) parameters in Table 4. Moreover, Table 5 display the training learning rates for models with different sizes.
Table 4: Architecture hyperparameters for GPT-2 series models (radford2019language).
@@
Table 5: Learning rates for GPT-2 experiments for different datasets.
@@ Appendix B Additional Experiments
We just display the curves for training and validation losses for small models on the FineWeb-Edu 100B in Figure 7. And we also display the curves for training and validation losses for the entire training process on the OpenWebText and FineWeb-Edu 100B datasets in Figures 8-9 and Figures 10-11, respectively. And the 0-shot and 2-shot evaluation results for these models on two datasets are shown in Tables 6-10 and Tables 11-16, respectively. The results of AdamW are taken from yuan2024mars. According to these figures and tables, it can be observed that MARS-M can achieve better performances than AdamW and Moonlight optimizers in most of these cases.
@@
Figure 8: The zoomed-in training loss of small-size (125M), medium-size (355M) and large-size (770M) models trained with different optimizers on the OpenWebText dataset.
Figure 9: The zoomed-in training and validation loss of large-size models (770M) trained with different optimizers on the OpenWebText dataset.
Figure 10: The zoomed-in training loss of medium-size (355M), large-size (770M) and XL-size (1.5B) models trained with different optimizers on the FineWeb-Edu 100B dataset.
Figure 11: The zoomed-in validation loss of medium-size (355M), large-size (770M) and XL-size (1.5B) models trained with different optimizers on the FineWeb-Edu 100B dataset.
@@
For terms for variance reduction, we have the following lemma from yuan2024mars:
Lemma C.1 (yuan2024mars).
@@
, and
$G_{t+1}:=(1-\beta_{t+1})\mathbb{E}\bigg\langle\bm{\Delta}_{t},\nabla f(\mathbf{X}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{X}_{t+1})\bigg\rangle+\beta_{t+1}\mathbb{E}\bigg\langle\bm{\Delta}_{t},\mathbf{M}_{t}-\nabla F(\mathbf{X}_{t})\bigg\rangle.$
We also need the following lemmas, which can be seen as an extension of Lemma C.3 in yuan2024mars.
@@
Here $\mathbf{O}_{t}=\mathbf{U}_{t}\mathbf{V}_{t}^{\top}$ is defined in (4.1).  [^3] In Algorithm 2, $\mathbf{O}_{t}$ is computed via a (truncated) Newton–Schulz iteration; for the purpose of this analysis, we treat $\mathbf{O}_{t}$ as the exact polar factor $\mathbf{U}_{t}\mathbf{V}_{t}^{\top}$ and ignore the approximation error.According to the definition of $\mathbf{O}_{t}$, we have:
@@
$\leq\frac{\rho_{t}}{16L^{2}}\cdot\Big(\frac{\beta_{t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big)\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\rho_{t}}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\frac{\rho_{t}\eta_{t}\sigma^{2}}{2L^{2}}-\frac{\rho_{t}}{16L^{2}\eta_{t}}\widetilde{M}_{t+1},$ (C.4)
@@
From Lemma C.3, we know that $\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\leq\eta_{t}^{1/2}$. Since $\rho_{t}^{2}=32L^{2}\eta_{t}$, we obtain [^5] As long as $\eta_{t}\leq 1/2$, $1\geq\beta_{t+1}=1-2\eta_{t}\geq 0$ is satisfied.
$\frac{\rho_{t}}{16L^{2}}\cdot\Big(\frac{\beta_{t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big)\leq\frac{\rho_{t}}{16L^{2}}\cdot\Big(\eta_{t}^{1/2}-2\Big)\leq\frac{\rho_{t}}{16L^{2}}\cdot\Big(1-2\Big)=-2\eta_{t}\rho_{t}^{-1}.$ (C.5)
@@
$I_{2}\leq-\frac{2\eta_{t}}{\rho_{t}}\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\rho_{t}}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\frac{\rho_{t}\eta_{t}\sigma^{2}}{2L^{2}}-\frac{\rho_{t}}{16L^{2}\eta_{t}}\widetilde{M}_{t+1}.$ (C.6)
@@
$\Phi_{t+1}-\Phi_{t}$ $\leq-\frac{\eta_{t}}{\rho_{t}}\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}-\eta_{t}\mathbb{E}\|\mathbf{M}_{t}\|_{F}+\frac{\rho_{t}}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\frac{\rho_{t}\eta_{t}\sigma^{2}}{2L^{2}}$
@@
$\leq\Phi_{1}-\Phi_{T+1}+\frac{\sigma^{2}}{2L^{2}}\sum_{t=1}^{T}\rho_{t}\eta_{t}+\sum_{t=1}^{T}\big(\frac{3\rho_{t}\eta_{t}n}{8}+\frac{L\eta_{t}^{2}n}{2}\big)-\sum_{t=1}^{T}\frac{\sqrt{2}}{4L\sqrt{\eta_{t}}}\widetilde{M}_{t+1}$